OpenStudy (anonymous):

Solve each equation for x, where a, b, c ER a) IaxI -b = c b) Ix-bI = c (the I are absolute value brackets)

OpenStudy (anonymous):

You have to split each equation into two cases.

OpenStudy (anonymous):

@ktm would you mind walking me through that?

OpenStudy (anonymous):

for a) we have |ax|-b = c then |ax| = c+b now we need to split it into two cases to get rid of the absolute value, so we have -ax = c+b and ax = c+b then you can solve those two equations for x